# Transport, Variational Inference and Diffusions: with Applications to Annealed Flows and Schrödinger Bridges

### v1

@@ E.4 Drift based EM
and we can conclude as in the proof of Proposition 4.3. ∎
Appendix F Numerical implementation
F.1 Discretisation and connection to DNFs (diffusion normalising flows)
Proposition F.1.
Letting $\Gamma_{0}=\Gamma_{T}=\mathrm{Leb}$ and $\gamma^{\pm}=0$, we have that the RND in (2.2) is given by
[…]
$\operatorname{ln}\left(\widehat{\frac{\mathrm{d}\overrightarrow{\mathbb{P}}^{\mu,a}}{\mathrm{d}\overleftarrow{\mathbb{P}}^{\nu,b}}}\right)({\bm{Y}})=-\operatorname{ln}\nu({\bm{Y}}_{T})+\sum_{i=0}^{K-1}\tfrac{1}{2\sigma^{2}\sqrt{t_{i+1}-t_{i+1}}}||{\bm{Y}}_{t_{i}}-{\bm{Y}}_{t_{i+1}}+b_{t_{i+1}}({\bm{Y}}_{t_{i+1}})(t_{i+1}-t_{i+1})||^{2}+\mathrm{const.},$
when using the Euler-Maruyama discretisation:
Proof.
The first part follows by direct computation.
Adding and subtracting ${\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}$ allows us to complete the square in each sum, resulting in:
$\operatorname{ln}\left(\frac{\mathrm{d}\overrightarrow{\mathbb{P}}^{\mu,a}}{\mathrm{d}\overleftarrow{\mathbb{P}}^{\nu,b}}\right)({\bm{Y}})$ $\approx\operatorname{ln}\mu({\bm{Y}}_{0})-\operatorname{ln}\nu({\bm{Y}}_{T})+\sum_{i=0}^{K-1}\tfrac{1}{2\sigma^{2}\sqrt{t_{i+1}-t_{i}}}||{\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}-a_{t_{i}}(t_{i+1}-t_{i})||^{2}$
$+\sum_{i=0}^{K-1}\tfrac{1}{2\sigma^{2}\sqrt{t_{i+1}-t_{i}}}||{\bm{Y}}_{t_{i}}-{\bm{Y}}_{t_{i+1}}+b_{t_{i+1}}(t_{i+1}-t_{i})||^{2}.$ (73)
Now notice that under the Euler-Maruyama discretisation $||{\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}-a_{t_{i}}(t_{i+1}-t_{i})||^{2}=\sigma^{2}||\xi||^{2}$ where $\xi\sim{\mathcal{N}}(0,I)$ does not depend on $a_{t}$ or $b_{t}$; in particular when using $D_{\mathrm{KL}}$ for the divergence we have that $\mathbb{E}_{\overrightarrow{\mathbb{P}}^{\mu,a}_{\mathrm{EM}}}||{\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}-a_{t_{i}}(t_{i+1}-t_{i})||^{2}=\sigma^{2}$ and thus:
$\operatorname{ln}\left(\widehat{\frac{\mathrm{d}\overrightarrow{\mathbb{P}}^{\mu,a}}{\mathrm{d}\overleftarrow{\mathbb{P}}^{\nu,b}}}\right)({\bm{Y}})\propto\operatorname{ln}\mu({\bm{Y}}_{0})-\operatorname{ln}\nu({\bm{Y}}_{T})+\sum_{i=0}^{K-1}\tfrac{1}{2\sigma^{2}\sqrt{t_{i+1}-t_{i}}}||{\bm{Y}}_{t_{i}}-{\bm{Y}}_{t_{i+1}}+b_{t_{i+1}}(t_{i+1}-t_{i})||^{2}.$ (74)
Notice that in expectation (for computing $D_{\mathrm{KL}}$), equation (74) matches equation (15) in Zhang and Chen [2021] and thus provides a theoretical backing to the objective used in Zhang and Chen [2021]. Resolving the term $\mathbb{E}_{\overrightarrow{\mathbb{P}}^{\mu,a}_{\mathrm{EM}}}||{\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}-a_{t_{i}}(t_{i+1}-t_{i})||^{2}$ analytically may offer a variance reduction similar to the analytic calculations in Sohl-Dickstein et al. [2015, Equation 14] and the Rao-Blackwelizations of $D_{\mathrm{KL}}$ in Ho et al. [2020].
Remark 17.
The time discretised RND in equation (F.1) can be expressed as the ratio of the transition densities corresponding to two discrete-time Markov chains $\mu({\bm{y}}_{0})q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})/p^{b}({\bm{y}}_{0:K-1}|{\bm{y}}_{K})\nu({\bm{y}}_{K})$ with ${\bm{y}}_{0:K}\sim q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})\mu({\bm{y}}_{0})$. As a result considering $\nu(x)=\hat{\nu}({\bm{x}})/Z$ and the IS estimator $\hat{Z}=p^{b}({\bm{y}}_{0:K-1}|{\bm{y}}_{K})\hat{\nu}({\bm{y}}_{K})/\mu({\bm{y}}_{0})q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})$ it follows that $\mathbb{E}_{q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})\mu({\bm{y}}_{0})}[\ln\hat{Z}]$ is an ELBO of $\hat{Z}$ (e.g. $\mathbb{E}_{q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})\mu({\bm{y}}_{0})}[\ln\hat{Z}]\leq\ln{Z}$).
Whilst superficially simple, Remark 17 guarantees that normalizing constant estimators arising from our discretisation do not overestimate the true normalizing constant. This result is beneficial in practice as it allows us to compare estimators possessing this property by selecting the one with the largest value. As highlighted in [Vargas et al., 2023a] many SDE discretisations can result in estimators that do not yield an ELBO: for example, the estimators used in [Berner et al., 2022] can result in overestimating the normalising constant. Note similar remarks have been established in the context of free energy computation and the Jarzynski equality see Stoltz et al. [2010, Remark 4.5].
F.2 Neural network parameterisations
Following Zhang and Chen [2021] and the recent success in score generative modelling we choose the following parameterisations:
$a_{t}({\bm{x}})$ $=f_{t}({\bm{x}})+\sigma^{2}\nabla\phi(t,{\bm{x}}),$ (75a)
where $s_{\theta}$ is a score network [Song et al., 2021, De Bortoli et al., 2021, Zhang and Chen, 2021] and $\phi(t,{\bm{x}})$ is a neural network potential. We adapt the architectures proposed in Onken et al. [2021], Koshizuka and Sato [2023] to general activation functions. Note that these architectures allow for fast computation of $\Delta\phi$ comparable to that of Hutchinson’s trace estimator [Grathwohl et al., 2019, Hutchinson, 1989].
[…]

### v6

@@ E.5 HJB-Regularizers
and we can conclude as in the proof of Proposition E.4. ∎
Appendix F CMCD Experiments
In this section, we will cover further details pertaining to our experimental setup.
F.1 ELBO experiments and Comparison to Geffner & Domke (2023)
We compare our underdamped and overdamped CMCD variants against 5 datasets from Geffner & Domke (2023), which we describe in further detail below.
  • log_sonar ($d=61$) and log_ionosphere ($d=35$) are Bayesian logistic regression models: $x\sim\mathcal{N}(0,\sigma_{w}^{2}I),y_{i}\sim\mathrm{Bernoulli}(\mathrm{sigmoid}(x^{\top}u_{i}))$ with posteriors conditioned on the sonar and ionosphere datasets respectively.
  • brownian ($d=32$) corresponds to the time discretisation of a Brownian motion:
    $\alpha_{\mathrm{inn}}$ $\sim\mathrm{LogNormal}(0,2),$
    $\alpha_{\mathrm{obs}}$ $\sim\mathrm{LogNormal}(0,2),$
    $x_{1}$ $\sim{\mathcal{N}}(0,\alpha_{\mathrm{inn}}),$
    $x_{i}$ $\sim{\mathcal{N}}(x_{i-1},\alpha_{\mathrm{inn}}),\quad i=2,\ldots 20,$
    $y_{i}$ $\sim{\mathcal{N}}(x_{i},\alpha_{\mathrm{obs}}),\quad i=1,\ldots 30.$
    inference is performed over the variables $\alpha_{\mathrm{inn}},\alpha_{\mathrm{obs}}$ and $\{x_{i}\}_{i=1}^{30}$ given the observations $\{y_{i}\}_{i=1}^{10}\cup\{y_{i}\}_{i=20}^{30}$.
  • lorenz ($d=90$) is the discretisation of a highly stiff 3-dimensional SDE that models atmospheric convection:
    $\begin{array}[]{rlrl}x_{1}&\sim\mathcal{N}(\text{ loc }=0,\text{ scale }=1)&\\ y_{1}&\sim\mathcal{N}(\text{ loc }=0,\text{ scale }=1)&\\ z_{1}&\sim\mathcal{N}(\text{ loc }=0,\text{ scale }=1)&&\\ x_{i}&\sim\mathcal{N}\left(\text{ loc }=10\left(y_{i-1}-x_{i-1}\right),\text{ scale }=\alpha_{\text{inn }}\right)&i=2,\ldots,30\\ y_{i}&\left.\sim\mathcal{N}\left(\text{ loc }=x_{i-1}\left(28-z_{i-1}\right)-y_{i-1}\right),\text{ scale }=\alpha_{\text{inn }}\right)&i=2,\ldots,30\\ z_{i}&\sim\mathcal{N}\left(\text{ loc }=x_{i-1}y_{i-1}-\frac{8}{3}z_{i-1},\text{ scale }=\alpha_{\text{inn }}\right)&i=2,\ldots,30,\\ o_{i}&\sim\mathcal{N}\left(\text{ loc }=x_{i},\text{ scale }=1\right)&i=2,\ldots,30\end{array}$
    where $\alpha_{\text{inn }}=0.1$ (determined by the discretization step-size used for the original SDE). The goal is to do inference over $x_{i},y_{i},z_{i}$ for $i=1,\ldots,30$, given observed values $o_{i}$ for $i\in\{1,\ldots,10\}\cup\{20,\ldots,30\}$.
  • seeds ($d=26$) is a random effect regression model trained on the seeds dataset:
    $\tau\sim\operatorname{Gamma}(0.01,0.01)$
    $a_{0}\sim\mathcal{N}(0,10)$
    $a_{1}\sim\mathcal{N}(0,10)$
    $a_{2}\sim\mathcal{N}(0,10)$
    $a_{12}\sim\mathcal{N}(0,10)$
    $b_{i}\sim\mathcal{N}\left(0,\frac{1}{\sqrt{\tau}}\right)$
    $i=1,\ldots,21$
    $\operatorname{logits}_{i}=a_{0}+a_{1}x_{i}+a_{2}y_{i}+a_{12}x_{i}y_{i}+b_{1}$
    $i=1,\ldots,21$
    $r_{i}\sim\operatorname{Binomial}\left(\operatorname{logits}_{i},N_{i}\right)$
    $i=1,\ldots,21\text{. }$
    The goal is to do inference over the variables $\tau,a_{0},a_{1},a_{2},a_{12}$ and $b_{i}$ for $i=1,\ldots,21$, given observed values for $x_{i},y_{i}$ and $N_{i}$.
For all target distributions, we follow the hyperparameter setup from Geffner & Domke (2023) from their code repository [^13] https://github.com/tomsons22/LDVI for all baseline methods (ULA, MCD, UHA, and LDVI) as well as our overdamped and underdamped variants. We first pretrain the source distribution to a mean-field Gaussian distribution trained for $150,000$ steps with ADAM and a learning rate of $10^{-2}$. We then train for $150000$ iterations with a batch size of $5$, tuning learning rate between $[10^{-5},10^{-4},10^{-3}]$ picking the best one based on mean ELBO after training. For all methods, during training the mean-field source distribution is continued to be trained, as well as the discretisation step size and $\epsilon=\delta t\sigma$. For the underdamped methods we also train the damping coefficient $\gamma$, and for methods involving a score network, i.e. MCD, LDVI, CMCD and CMCD (UD), we train the networks which are chosen to be fully-connected residual networks with layer sizes of $[20,20]$. In order to report the mean ELBO after training, we obtain 500 samples with 30 seeds and report an averaged value over them.
F.2 $\ln Z$, sample quality experiments and comparison to (Zhang & Chen, 2022; Vargas et al., 2023a)
Furthermore, we also include comparisons to a large-dimensional target distribution and two standard distributions with known $\ln Z$ replicated from Vargas et al. (2023a), which we summarise below.
  • lgcp ($d=1600$) is a high-dimensional Log Gaussian Cox process popular in spatial statistics (Møller et al., 1998). Using a $d=M\times M=1600$ grid, we obtain the unnormalised target density $\mathcal{N}(x;\mu,K)\prod_{i\in[1:M]^{2}}\exp\left(x_{i}y_{i}-a\exp\left(x_{i}\right)\right)$.
  • funnel ($d=10$) is a challenging distribution given by   $\pi_{T}(x_{1:10}=\mathcal{N}(x_{1};0,\sigma_{f}^{2})\mathcal{N}(x_{2:10};0,\exp(x_{1})I)$, with $\sigma_{f}^{2}=9$ (Neal, 2003).
  • gmm ($d=2$) is a two-dimensional Gaussian mixture model with three modes, given by the following target distribution
    $\pi_{T}(x)$ $=\frac{1}{3}\mathcal{N}\left(x;\begin{bmatrix}3\\ 0\end{bmatrix},\begin{bmatrix}0.7&0\\ 0&0.05\end{bmatrix}\right)+\frac{1}{3}\mathcal{N}\left(x;\begin{bmatrix}-2.5\\ 0\end{bmatrix},\begin{bmatrix}0.7&0\\ 0&0.05\end{bmatrix}\right)$
    $\ \ +\mathcal{N}\left(x;\begin{bmatrix}2\\ 3\end{bmatrix},\begin{bmatrix}1&0.95\\ 0.95&1\end{bmatrix}\right)$
For these target distributions, we follow the hyperparameter setup from Vargas et al. (2023a) from their code repository [^14] https://github.com/franciscovargas/denoising_diffusion_samplersfor the baseline methods of DDS and PIS, and replicate them as closely as possible for CMCD. Unlike the previous, we don’t pretrain the mean-field Gaussian source distribution $\mathcal{N}(0,\sigma_{\text{init}}^{2}I)$. We select the optimal learning rate in $[10^{-3},10^{-4},10^{-5}]$, the optimal standard deviation of the source distribution $\sigma_{\text{init}}$ in $[1,2,3,4,5]$ and the optimal $\alpha$ in $[0.1,0.5,1,1.5,2]$. Instead of training $\epsilon=\delta t\sigma$, we sweep over an optimal value in $[10^{-2},10^{-1},1]$. The models are trained with a batch size of $300$ for $11000$ steps, where we keep the source distribution parameters fixed, as well as $\epsilon$. For evaluation, we use 30 seeds with a batch size of 2000, and report average performance over the seeds. DDS and PIS use a 128-dimensional positional embedding, along with an additional network for the time parameters, however MCMD uses a regular score network. In order to make exact comparisons, we select differing network architecture sizes that result in an equivalent number of parameters for funnel and gmm. For lgcp, due to the high dimensionality of the dataset, we choose a small network for CMCD. We summarise these below. For gmm and funnel, it is possible to sample from the target distribution, and we report an OT-regularised distance ($\mathcal{W}_{2}^{\gamma}$) with a regularisation $\gamma=10^{-2}$. Similar to the mean ELBO, we draw 2000 samples from the models and the targets, and average $\mathcal{W}_{2}^{\gamma}$ over 30 seeds. We use the Python Optimal Transport [^15] https://pythonot.github.io/ library’s default implementation of entropy-regularised distance. Results for comparisons to DNF can be found in Table 5.
Figure 2: Architecture from (Geffner & Domke, 2023) used across experiments for our CMCD drift network. Softplus activations are used.
Table 1: Network Sizes for comparison. Note that CMCD has less parameters for the despite the Funnel target despite the larger drift due to the PIS and DDS networks having an additional grad network.
GMM LGCP Funnel
DDS $[10,10]$ $[64,64]$ $[64,64]$
PIS $[10,10]$ $[64,64]$ $[64,64]$
CMCD $[38,38]$ $[64,64]$ $[110,110]$
F.3 Comparisons with the log-variance loss - Mode collapse failure mode
Here, we report performance using the log-variance divergence-based loss (Nüsken & Richter, 2021) introduced at the end of Section 3.2,
$\mathcal{L}^{\mathrm{CMCD}}_{\mathrm{Var}}(\phi)\!\approx\!\mathrm{Var}\!\left[\!\ln\!\frac{{\pi}_{0}({\bm{Y}}_{0})}{\hat{\pi}({\bm{Y}}_{T})}\!\!\prod_{k=0}^{K-1}\!\!\frac{{\mathcal{N}}({\bm{Y}}_{t_{k+1}}|{\bm{Y}}_{t_{k}}+(\nabla\ln\pi_{t_{k}}+\nabla\ln\phi_{t_{k}})({\bm{Y}}_{t_{k}})\Delta t_{k},2\sigma^{2}\Delta t_{k})}{{\mathcal{N}}\!({\bm{Y}}_{t_{k}}|{\bm{Y}}_{t_{k+1}}\!\!\!+\!(\nabla\!\ln\pi_{t_{k+1}}\!\!\!-\!\nabla\!\ln\phi_{t_{k+1}})({\bm{Y}}_{t_{k+1}})\Delta t_{k},2\sigma^{2}\Delta t_{k})}\!\right]\!,$ (24)
A careful reader will note this loss simply consists of replacing the expectation in the KL loss with a variance. A major computational advantage of this loss is that the measure that the expectations are taken with respect to can be any measure and is not restricted to the forward or backward SDEs like in KL (Richter & Berner, 2024; Richter et al., 2020; Nüsken & Richter, 2021), this allows us to detach the samples and thus accommodating for a much more computational objective.
which we find performs quite well compared to our default loss function, especially for multimodal target distributions. We consider the very multi-modal mixture of Gaussian target distribution from Midgley et al. (2022), and report the ELBO and $\ln Z$ numbers in the table below. For this experiment, we use a batch size of 2000 and train neural networks with a size $[130,130]$ for $150k$ iterations.
Table 2: ELBO and $\ln Z$ on 40-GMM
ELBO $\ln Z$ $\mathcal{W}_{2}$
log-variance loss -1.279 $\pm$ 0.096 -0.065 $\pm$ $0.101$ 0.0143 $\pm$ 0.001
KL loss -2.286 $\pm$ 0.1109 -0.244 $\pm$ 0.3309 0.0441 $\pm$ 0.012
Figure 3: (left) 2000 samples drawn from the CMCD algorithm trained with the default loss function, and (right) 2000 samples drawn from the algorithm trained with the log-variance divergence-based loss. We can see that the default loss function misses many modes in the target distribution, whereas the log-variance loss has not missed any modes. We report final results after sweeping over $\Delta_{t_{k}}$ and learning rates for both methods, picking the one with the lowest training loss. We highlight that concurrent work by Richter & Berner (2024) explores the log variance divergence in more detail and proposes an akin general framework for diffusion-based sampling.
Figure 4: Plots showing training loss curves for the log-variance loss and the default loss for different values of $\Delta_{t_{k}}$. We find that a low value of $\Delta_{t_{k}}=0.1$ is needed in order to obtain a low training loss for the default loss, whereas the log-variance loss is much more robust to different values of $\Delta_{t_{k}}$. The x-axis reports an evaluation every 150 steps of training
F.4 Specification and Tuning: SMC, AFT, and NF-VI
We adopted the implementations [^16] https://github.com/google-deepmind/annealed_flow_transport provided by the studies in Arbel et al. (2021); Matthews et al. (2022) and initialized them with default hyperparameters before fine-tuning.
Sequential Monte Carlo (SMC). For SMC, we utilized 2000 particles sampled from a zero-mean unit Gaussian distribution, implementing re-sampling if the effective sample size (ESS) fell below 0.3. We employed Hamiltonian Monte-Carlo (HMC) for particle mutation, executing one Markov Chain Monte Carlo (MCMC) step after each annealing step. The number of leapfrog steps was fixed at 10, and an extensive grid search over different step sizes was conducted, consistent with Arbel et al. (2021). This search spanned four different step sizes, contingent on the temperature, resulting in a grid search over 256 parameters. The finalized values are presented in Table 3. For SMC, $K$ is defined by the number of temperatures.
Furthermore, we report the results for SMC in Table 3 for the tuned hyperparameters used for each target. Please note that we were able to obtain similar $\ln Z$ values as in Vargas et al. (2023a) suggesting SMC was well-tuned. Finally, for each result, we report the mean and standard deviations across 30 different seeds, results can be seen in Table 3.
Annealed Flow Transport Monte Carlo (AFT). We maintained a similar setup to SMC, with a few adjustments: using 500 particles for training and 2000 for evaluation to accommodate the added complexity from the normalizing flows. We also decreased the number of temperatures and increased the number of MCMC steps to mitigate memory requirements from the flows. $K$ is defined as the number of temperatures $\times$ MCMC steps, with the latter fixed at 4, resulting in a maximum of 64 flows trained simultaneously. Inverse autoregressive flows (IAFs) were employed in all experiments except for lgcp, using a neural network with one hidden layer whose dimension matches the problem’s dimensionality. For lgcp, a diagonal affine flow was used due to memory constraints arising from the high dimensionality. AFT flows were trained for 300 iterations until convergence.
Variational Inference with Normalizing Flows (VI-NF). We utilized the same flows as for AFT. In this case, $K$ denotes the number of flows to stack. The flows were trained over a total of 2000 iterations with a batch size of 500. For some targets
Table 3: Tuned MCMC Step Sizes.
gmm lgcp lorenz brownian log_sonar log_ionosphere seeds funnel
$\Delta t$ $[0.5,0.5,0.5,0.3]$ $[0.3,0.3,0.2,0.2]$ $[0.01,0.01,0.008,0.01]$ $[0.2,0.2,0.05,0.05]$ $[0.2,0.05,0.2,0.2]$ $[0.1,0.2,0.2,0.2]$ $[0.2,0.1,0.05,0.01]$ $[0.05,0.2,0.2,0.05]$
Table 4: SMC Results. ELBO and $\ln Z$ values for a different number of steps $K$ and experiments.
$\ln Z$ gmm lgcp lorenz brownian log_sonar log_ionosphere seeds funnel
$K=8$ $-0.536\pm 0.042$ $-364.074\pm 7.797$ $-87502.352\pm 4004.495$ $-63.32\pm 8.016$ $-178.589\pm 2.784$ $-204.594\pm 3.049$ $-108.676\pm 1.221$ $-1.013\pm 0.116$
$K=16$ $-0.255\pm 0.034$ $-135.207\pm 4.665$ $-42148.287\pm 1047.478$ $-28.714\pm 3.71$ $-137.691\pm 1.656$ $-149.107\pm 1.088$ $-88.068\pm 0.467$ $-0.65\pm 0.1$
$K=32$ $-0.119\pm 0.017$ $86.106\pm 5.989$ $-19288.267\pm 834.52$ $-12.23\pm 2.212$ $-120.557\pm 0.613$ $-127.964\pm 0.394$ $-79.89\pm 0.273$ $-0.408\pm 0.17$
$K=64$ $-0.059\pm 0.015$ $269.566\pm 7.832$ $-8894.525\pm 119.723$ $-4.76\pm 1.042$ $-113.835\pm 0.167$ $-118.812\pm 0.192$ $-76.275\pm 0.189$ $-0.359\pm 0.087$
$K=128$ $-0.029\pm 0.009$ $390.33\pm 5.427$ $-5419.678\pm 90.362$ $-1.675\pm 0.442$ $-110.901\pm 0.094$ $-114.827\pm 0.307$ $-74.774\pm 0.097$ $-0.255\pm 0.108$
$K=256$ $-0.013\pm 0.006$ $477.162\pm 4.998$ $-3745.218\pm 68.342$ $-0.131\pm 0.22$ $-109.562\pm 0.072$ $-113.123\pm 0.172$ $-74.049\pm 0.088$ $-0.211\pm 0.074$
ELBO
$K=8$ $0.002\pm 0.066$ $-236.087\pm 9.623$ $-56122.917\pm 5402.094$ $-10.147\pm 3.427$ $-117.499\pm 4.049$ $-123.772\pm 3.689$ $-75.183\pm 1.447$ $-0.417\pm 0.236$
$K=16$ $-0.003\pm 0.037$ $-23.219\pm 7.756$ $-27397.2\pm 1987.523$ $-3.924\pm 2.114$ $-110.707\pm 1.823$ $-113.476\pm 1.361$ $-73.524\pm 0.543$ $-0.322\pm 0.184$
$K=32$ $0.003\pm 0.018$ $174.797\pm 7.241$ $-12110.983\pm 1204.2$ $-0.426\pm 1.415$ $-108.574\pm 0.547$ $-112.048\pm 0.519$ $-73.459\pm 0.29$ $-0.215\pm 0.222$
$K=64$ $0.001\pm 0.015$ $332.187\pm 9.025$ $-5360.819\pm 306.407$ $0.884\pm 0.778$ $-108.424\pm 0.154$ $-111.715\pm 0.184$ $-73.375\pm 0.214$ $-0.267\pm 0.101$
$K=128$ $0.001\pm 0.009$ $430.838\pm 6.441$ $-3624.167\pm 168.119$ $1.008\pm 0.27$ $-108.395\pm 0.087$ $-111.603\pm 0.298$ $-73.436\pm 0.095$ $-0.2\pm 0.124$
$K=256$ $0.002\pm 0.006$ $453.395\pm 4.43$ $-2811.161\pm 106.68$ $1.142\pm 0.125$ $-108.368\pm 0.071$ $-111.611\pm 0.171$ $-73.413\pm 0.087$ $-0.181\pm 0.081$
F.5 Further Ablation with NF-style methods and AFT
We further run both flow models (AFT and NFVI) on all possible target distributions (subject to OOM errors). Results can be found in Table 6.
Table 5: $\ln Z$ comparison. $\ln Z$ values for a different number of steps $K$, experiments and methods. Not all methods could be evaluated on every $K$/experiment combination due to numerical instabilities or out-of-memory (OOM) problems.
Dataset Method $K=8$ $K=16$ $K=32$ $K=64$ $K=128$ $K=256$
funnel CMCD -0.3037 $\pm$ 0.1507 -0.223 $\pm$ 0.1041 -0.1805 $\pm$ 0.0773 -0.1085 $\pm$ 0.1143 -0.0573 $\pm$ 0.0444 -0.01928 $\pm$ 0.0641
log_ionosphere VI-DNF -0.3768 $\pm$ 0.2157 -0.3517 $\pm$ 0.1627 -0.2919 $\pm$ 0.0999 -0.6941 $\pm$ 0.6841 -0.1947 $\pm$ 0.1325 -0.2124 $\pm$ 0.0637
VI-NF -0.206$\pm$ 0.079 -0.206$\pm$ 0.082 -0.206$\pm$ 0.087 -0.194$\pm$ 0.101 -0.182$\pm$ 0.097 -0.197$\pm$ 0.099
AFT -0.875$\pm$ 0.543 -0.395$\pm$ 0.351 -0.348$\pm$ 0.192 -0.271$\pm$ 0.227 -0.235$\pm$ 0.139 -0.196$\pm$ 0.111
gmm CMCD -0.1358 $\pm$ 0.0839 -0.01331 $\pm$ 0.1292 0.0095 $\pm$ 0.0495 0.00736 $\pm$ 0.0477 -0.0004 $\pm$ 0.0368 -0.0081 $\pm$ 0.0520
VI-DNF -0.3676 $\pm$ 0.6314 -0.258 $\pm$ 0.412 -0.4983 $\pm$ 0.3878 -0.4449 $\pm$ 0.5379 -0.4652 $\pm$ 0.3223 -0.204 $\pm$ 0.6381
VI-NF -0.355$\pm$ 0.698 -0.455$\pm$ 0.258 -0.064$\pm$ 0.138 -0.054$\pm$ 0.15 -0.066$\pm$ 0.188 -0.045$\pm$ 0.177
AFT -0.336$\pm$ 0.372 -0.006$\pm$ 0.082 0.02$\pm$ 0.068 -0.016$\pm$ 0.042 -0.003$\pm$ 0.029 0.001$\pm$ 0.026
lgcp CMCD 491.059 $\pm$ 3.553 498.147 $\pm$ 2.624 502.705 $\pm$ 2.482 506.045 $\pm$ 1.761 508.165 $\pm$ 1.553 509.43 $\pm$ 1.242
VI-DNF 424.733 $\pm$ 5.858 424.719 $\pm$ 5.855 424.714 $\pm$ 5.861 424.719 $\pm$ 5.860 424.7 $\pm$ 5.869 424.705 $\pm$ 5.896
AFT 126.651$\pm$ 5.764 344.145$\pm$ 23.95 191.613$\pm$ 173.873 420.259$\pm$ 91.43 480.126$\pm$ 33.059 491.028$\pm$ 8.057
Table 6: ELBO comparison. ELBO values for a different number of steps $K$, experiments and methods. Not all methods could be evaluated on every $K$/experiment combination due to numerical instabilities or out-of-memory (OOM) problems.
Dataset Method $K=8$ $K=16$ $K=32$ $K=64$ $K=128$ $K=256$
seeds CMCD $-74.501\pm 0.049$ $-74.327\pm 0.065$ $-74.142\pm 0.05$ $-73.967\pm 0.038$ $-73.8\pm 0.032$ $-73.684\pm 0.033$
VI-NF $-73.563\pm 0.013$ $-73.547\pm 0.012$ $-73.574\pm 0.012$ $-73.58\pm 0.014$ $-73.621\pm 0.014$ $-73.675\pm 0.014$
AFT $-147.457\pm 24.808$ $-116.134\pm 8.157$ $-99.032\pm 6.321$ $-87.436\pm 1.53$ $-79.847\pm 0.419$ $-76.364\pm 0.188$
CRAFT $-146.973\pm 1.531$ $-94.2\pm 0.505$ $-80.985\pm 0.344$ $-76.555\pm 0.175$ $-74.979\pm 0.143$ $-74.225\pm 0.133$
log_ionosphere CMCD $-113.211\pm 0.089$ $-112.643\pm 0.062$ $-112.643\pm 0.062$ $-112.22\pm 0.046$ $-111.98\pm 0.04$ $-111.925\pm 0.046$
VI-NF $-111.903\pm 0.022$ $-111.902\pm 0.022$ $-111.892\pm 0.017$ $-111.881\pm 0.017$ OOM OOM
AFT $-168.174\pm 21.249$ $-138.733\pm 8.374$ $-123.013\pm 3.771$ $-118.644\pm 0.891$ $-116.497\pm 0.495$ $-114.905\pm 0.781$
log_sonar CMCD $-112.274\pm 0.124$ $-110.904\pm 0.111$ $-110.459\pm 0.106$ $-109.503\pm 0.075$ $-109.608\pm 0.066$ $-109.25\pm 0.052$
VI-NF $-109.353\pm 0.035$ $-109.346\pm 0.031$ $-109.441\pm 0.035$ $-109.94\pm 0.044$ $-109.711\pm 0.039$ OOM
AFT $-203.249\pm 12.506$ $-148.357\pm 8.096$ $-129.772\pm 3.057$ $-121.653\pm 2.505$ $-114.911\pm 0.331$ $-112.021\pm 0.182$
lgcp CMCD $469.475\pm 0.259$ $479.246\pm 0.237$ $486.739\pm 0.249$ $492.745\pm 0.239$ $497.074\pm 0.267$ $499.708\pm 0.236$
AFT $75.896\pm 0.863$ $265.005\pm 34.254$ $62.898\pm 200.991$ $340.687\pm 126.853$ $417.916\pm 50.35$ $424.705\pm 12.416$
lorenz CMCD $-1180.797\pm 0.184$ $-1180.797\pm 0.184$ $-1176.514\pm 0.154$ $-1174.309\pm 0.148$ $-1172.453\pm 0.153$ $-1170.826\pm 0.15$
VI-NF $-1499.102\pm 0.84$ $-1471.798\pm 0.582$ $-1439.648\pm 0.274$ $-1433.536\pm 0.316$ OOM OOM
brownian CMCD $-0.753\pm 0.075$ $-0.209\pm 0.059$ $0.153\pm 0.045$ $0.376\pm 0.038$ $0.578\pm 0.046$ $0.722\pm 0.032$
VI-NF $0.733\pm 0.019$ $0.797\pm 0.018$ $0.816\pm 0.018$ OOM OOM OOM
F.6 Wallclock times for $\operatorname{ln}Z$ calculation
In order to calculate the average wall-clock time for $\operatorname{ln}Z$ calculation, we calculate the time it takes to draw $30$ seeds of $2000$ samples each from the methods below, and use these samples to calculate the mean and standard deviation of $\operatorname{ln}Z$ across $30$ seeds.
Method Average Time (s) Min Time (s) Max Time (s)
CMCD (OD) 9.665 5.592 21.475
ULA 9.204 4.673 20.721
UHA 9.427 5.588 20.263
MCD 9.204 4.673 20.721
Table 7: Wallclock times for evaluation in seconds.
Mode ULA MCD CMCD DDS / PIS Sampling $\mathcal{O}(K\cdot(d+G(d)))$ $\mathcal{O}(K\cdot(d+G(d)))$ $\mathcal{O}(K\cdot(d+G(d)+N(d)))$ $\mathcal{O}(K\cdot(N_{2}+G(d)+N_{1}(d)))$ ELBO $\mathcal{O}(K\cdot(d+G(d)))$ $\mathcal{O}(K\cdot(d+G(d)+N(d)))$ $\mathcal{O}(K\cdot(d+G(d)+N(d)))$ $\mathcal{O}(K\cdot(N_{2}+G(d)+N_{1}(d)))$
Table 8: Sampling and loss calculation complexity across SDE based methods, $K$ represents the number of integration steps, $G(d)$ represents the cost of evaluating the score of the target and $N(d)$ for evaluating the drift/score networks both quantities are dimension dependant. PIS and DDS have an additional grad network cost $N_{2}$ which is dimension independant.
F.7 Training time comparisons to SMC
In this section, we explore a total time comparison between our approach CMCD and SMC.
As both methods are quite inherently different it is not immediately obvious how to carry out an insightful comparison. In order to do so we chose the LGCP which is our most numerically intense target and we phrase the following question:
“For how long do we have to train CMCD to outperform the best-run SMC”
For this, we look at our Figure 1 pane c) and we can see that at $K=8$ CMCD already outperforms SMC at $K=256$ with 2000 particles. So we choose these two approaches to compare to. In Table 9 is a brief comparison of total time calculations, note we have included tuning time for SMC which is akin to our training time as without tuning SMCs hyperparameters ELBOs and ln Z estimations were much worse. We can observe that the total runtime for training and sampling CMCD to reach a better $\ln Z$ value does not exceed the time required to tune SMC.
Method Train + Sample Time (min) ln Z ELBO
CMCD $33.12\pm 0.12$ $491.059\pm 3.553$ $469.475\pm 0.2589$
SMC $62.62\pm 0.10$ $477.162\pm 4.998$ $453.395\pm 4.4300$
Table 9: Training+ Tuning + Sampling time comparisons for CMCD and SMC at comparable ln Z estimates.
Appendix G Regularised IPF-type Experiments
[…]